aiaa – a L^{A}TEX Class and BibTEX Style for AIAA^{∗} Conference Papers and Journal Submission/Simulation^{†}

Bil Kleb^{‡}

1997/02/20

Abstract

This document describes the aiaa distribution which is centered around a modification of the standard L^{A}TEX article class, article.cls. The new class produces AIAA-conformant^{1} conference papers and journal submittals—it will even simulate the typesetting of journal articles and notes for lengthdetermination purposes. This distribution also contains a (mostly) AIAAcompliant bibliographic style sheet, a sample document, and some other test equipment.

Contents

1 Introduction 2

2 Userguide 2 2.1 Requirements . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.2 Setup . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 2.3 Usage . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 2.4 General Commands . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

2.4.1 New Behavior from Old Commands . . . . . . . . . . . . . 3 2.4.2 New Commands . . . . . . . . . . . . . . . . . . . . . . . . 4

2.5 Layout-specific Commands . . . . . . . . . . . . . . . . . . . . . . . 4 2.5.1 Header and Footer Information . . . . . . . . . . . . . . . . 5 2.5.2 Cover-page Information . . . . . . . . . . . . . . . . . . . . 5 2.5.3 Copyright and Other Document Notices . . . . . . . . . . . 5

3 Known Problems 6

4 Acknowledgements 7

5 Sending a Bug Report 7

∗The American Institute of Aeronautics and Astronautics.

†This document describes aiaa version 0.9, and was last updated 1997/02/20.

‡Research Scientist, NASA Langley Research Center, Hampton, Virginia.

1Note this distribution is not a product of, nor endorsed by, the AIAA.

1 Introduction

The aiaa distribution consists of a L^{A}TEX class and various other files which simplifies the task of producing an AIAA conference paper and the subsequent journal submission For instance, with a simple, one-word option in the beginning of the document, a manuscript can take the form of a two-column conference paper replete with cover-page or a double-spaced journal submission manuscript with figures and tables at the end, including the required caption lists.

2 Userguide

I apologize for the sparseness of this documentation; but, hey, this is not in my job description. ;) Hopefully the sample document smpaiaa.tex provides suitable documentation through example.

2.1 Requirements

The aiaa distribution was developed using L^{A}TEX 2ε of 1995/12/01, patch level 1, running TEX 3.14159 and dvips 5.58f—the teTEX distribution 0.3.3. The aiaa distribution has also been demonstrated on a Mac running the OzTEX TEX distribution.

The aiaa distribution works best with dvips and MakeTeXPK installed, although not formally needed. Without these, the output may not be as good as expected, affecting mainly the appearance of the dropped capitals which can be incorporated in the first paragraph of the typeset text.

2.2 Setup

If you have not already run aiaa.ins through L^{A}TEX, do so know. The docscript package will rip the code bits out of aiaa.dtx and save them in several

files. Move the files aiaa.cls, aiaa9pt.sty, aiaaenf.cfg, and aiaalgo.eps to

a directory/folder searched by TEX^{2} and the file aiaa.bst to a directory/folder searched by BibTEX.^{3}

The aiaa class depends on having access to a number of packages.^{4} If your local site does not have all the packages necessary, they can be obtained from your nearest Comprehensive TeX Archive Network (CTAN) site. In fact, chances are that this is where you found this distribution. Details on how to obtain packages from a CTAN site are available at http://www.tug.org/ or various L^{A}TEX reference books [1, 2]. Especially helpful in locating various L^{A}TEX packages is the Catalogue.html web page found in the help/Catalogue directory of CTAN.

2For a Unix teTEX installation, a privileged user could put these files in a directory named something like /usr/local/teTeX/texmf/tex/latex/aiaa for the entire site to use, remembering to run texhash to re-configure teTEX to include the new directory; or, a lowly user could make their own directory, e.g., ∼/tex/inputs, put the files in there, and set the environment variable TEXINPUTS via ‘setenv TEXINPUTS ∼/tex/inputs:’. The colon represents the system search path so, in this case, the user files take precedence. On a Mac or PC installation put these files in a folder named something like TeX-inputs.

3Similar to preceding footnote, only on Unix use the environment variable BSTINPUTS for the bibliographic style file and BIBINPUTS for the bibliographic database; for Mac’s, use the BibTeX-inputs folder, failing that try using the TeX-inputs folder.

4When you process a file with L^{A}TEX, it will let you which packages it is missing. The less common ones are dropping and caption2 while the other required packages: lastpage, setspace, endfloat, overcite, graphicx, and fancyhdr are usually bundled with most L^{A}TEX distributions.

2.3 Usage

The aiaa class is envoked by including

  \documentclass[options]{aiaa}

at the beginning of your document. The package recognizes the following options:

submit, paper, article, note, and cover. The paper option is the default and

the submit option overrides any other option. The cover option produces a coverpage for conference papers or simulated journal reprints. In addition, any options that the standard L^{A}TEX article class accepts can be also inserted, e.g., draft.^{5} Note, when using the note option, your title and author information may need to be modified with extra line breaks (\\) since this information is no longer allowed to span both columns. Other than this, the document is written just like one were using the standard L^{A}TEX article document class; and thus, I don’t have to write much more on usage since it has already been documented by others in various L^{A}TEX books [3, 1, 2]. However, some of the old commands have different behaviors and there are a few new commands designed to make life a little brighter; these are discussed in the following sections.

2.4 General Commands

Several standard L^{A}TEX commands have been modified to behave differently under the new class. In addition, several new commands have been introduced to ease document preparation. Both types are discussed in the following subsections.

2.4.1 New Behavior from Old Commands

\abstract The \abstract command has been redefined within the aiaa class to behave as part of the \maketitle sequence. Simply load \abstract as you would \title or \author, and if you have not selected the options submit or note, the text loaded into the \abstract command will be typeset, indented and centered, underneath the title/author section.

\date The aiaa class automatically nulls the \date command used by \maketitle. Standard L^{A}TEX behavior of \maketitle is to typeset the current date as part of the title section if one is not given. So, normal one issues the command: \date{}. When using the aiaa class, this is not necessary; this command has already been issued.

\and Suffice it to say that \maketitle and \and have also been hacked. \maketitle One no longer has to use the starred versions of these commands to defeat

\section
\subsection
\subsubsection
\paragraph
\subparagraph

the numbering of sections. The counter secnumdepth has been set to −2 via, \setcounter{secnumdepth}{-2} so that even the unstarred versions of the sectioning commands never produce a number.^{6} Of course you can defeat this by changing the counter back to its L^{A}TEX article class default value of 3 which will number sections down to subsections.

5The draft option replaces figures with a properly-sized, blank, framed box.

6The secnumdepth counter controls how many nesting levels of section numbers should be produced.

In addition, the fonts, sizes, and positions normally produced by these commands have been modified to make the output similar to a typeset journal article. In other words, we take advantage of the fact that you are not using a typewriter, e.g., AIAA-suggested underlining for section names is replaced by boldface, etc.

2.4.2 New Commands

\dropword The command \dropword is used for the first word of the introduction, and for any option other than submit, will produce a dropped capital for the beginning of the paragraph. Use is simply:

  \dropword First words of the introduction, etc.

Note: this command also capitalizes the remaining portion of the first word. This macro relies on the presence of the dropping package by Mats Dahlgren [4].

\Figure The commands \Figure and \Reference automatically modulate between \Reference “Figure” or “Fig.” and “Reference” or “Ref.” depending on their location within

a sentence.^{7} These were created because the AIAA requires these words to be abbreviated when they do not begin a sentence.^{8} So, instead of

  Figure~\ref{f:fig1} is loud and Fig.~\ref{f:fig2} isn't.

you would type

  \Figure{f:fig1} is loud and \Figure{f:fig2} isn't.

Similarly for \Reference.

\incfig The \incfig command is used for including figures via the graphicx package of David Carlisle [5]. It is typically used for including Encapsulated Postscript files (*.eps) within the figure environment. Thus, to include a figure named figa.eps residing in a figs subdirectory, one would use:

   and \Figure{f:figurea} shows that things are purple.
   \begin{figure}
      \incfig{figs/figa}
      \caption{This is the caption for figure a.}
      \label{f:figurea}
   \end{figure}
   \Figure{f:figurea} also shows that people are green.

For more information regarding the inclusion of external graphic files, see the file epslatex.ps in the info directory on CTAN.

2.5 Layout-specific Commands

The following commands are used to load information for other commands that produce appropriate headers, footers, cover-page items, and document notices— e.g., copyright conditions. All of these commands are normally set in the preamble^{9} of your document (similar to \author and \title). For a more modular, and perhaps cleaner approach, you could place all of it in a file, preamble.tex, and

use \input{preamble} to include it.

7Beware, in bizarre instances these might fail, although they haven’t to date.

8So, what? … we save two whole characters and produce an abomination in the process? Oh well. You can defeat this abbreviation business and still use the short-hand for the cross referencing by issuing the commands \renewcommand{\Figure}{Figure∼\ref{#1}} and \renewcommand {\figurename} {Figure}.

9The preamble is defined as anywhere between the \documentclass{} and \begin{document} commands.

2.5.1 Header and Footer Information

\SubmitName The commands \SubmitName, \PaperNumber, \ArticleIssue, \ArticleHeader,
\PaperNumber
\ArticleIssue
\ArticleHeader
\NoteHeader

and \NoteHeader are used to put appropriate items in the header and footer of each page, e.g.,

   \SubmitName{Kleb}
   \PaperNumber{96--0825}
   \ArticleIssue{Vol.~32, No.~6, November--December 1995}% first page
   \ArticleHeader{Kleb et al: Pitch-Over Maneuver}% subsequent pages
   \NoteHeader{J.Spacecraft, Vol.~32, No.~6: Engineering Notes}

The command \SubmitName is used to mark the main author’s name on all pages for the journal manuscript submission option: submit. The \PaperNumber is used to include the paper number for AIAA conference papers. The commands

\ArticleIssue, \ArticleHeader, and \NoteHeader are used to create the ap-

propriate headers when simulating a journal article or note. The contents of \ArticleIssue appear on the first page and the contents of \ArticleHeader appear on subsequent pages while for journal note simulations the contents of \NoteHeader is used for all pages.

2.5.2 Cover-page Information

\CoverFigure The commands \CoverFigure, \Conference, \JournalIssue, and \JournalName
\Conference
\JournalName
\JournalIssue

provide information for producing cover-pages.

   \CoverFigure{tstfig.eps}
   \Conference{{\bfseries 31st AIAA Aerospace Sciences \\
                           Meeting and Exhibit} \\
               January 6--9, 1997/Reno, NV}
   \JournalName{Journal of Spacecraft and Rockets}
   \JournalIssue{Volume 32, Number 6, Pages 715--717}

The commands \CoverFigure and \Conference define a representative figure (optional) and conference name/date/location to be used on the cover-page of a conference paper, while the commands \JournalName and \JournalIssue are for the cover-page produced for journal article or note reprint simulations.

2.5.3 Copyright and Other Document Notices

\PaperNotice A footnote describing the copyright conditions and other information about the

\JournalNotice document are incorporated via the \PaperNotice and \JournalNotice com-

mands. These normally include one of the the copyright series of commands:

\copyrightA, \copyrightB, \copyrightC, or \copyrightD, described below. To

use, simply include something like the following in the your document’s preamble:

   \PaperNotice{\copyrightA{1996}}
   \JournalNotice{Presented as Paper 96--0825 at the AIAA 34th
                  Aerospace Sciences Meeting, Reno,~NV,
                  Jan.~15--18,~1996; received Feb.~15,~1996;
                  revision received Nov.~25,~1996. \copyrightC}

\copyrightA The copyright commands will expand to one of the standard AIAA forms: A,

\copyrightB
\copyrightC
\copyrightD

B, C, or D. They each have different arguments—or no arguments—depending on the requirements:

\copyrightA{year}
\copyrightB{year}{fullname}
\copyrightC
\copyrightD{year}

See your AIAA copyright instructions for which form to use.

3 Known Problems

• The bibliographic style sheet aiaa.bst isn’t fully tested; and thus, you may need to fiddle with your .bbl file for your final copy, i.e., edit file.bbl after running a L^{A}TEX, BibTEX, L^{A}TEX sequence, but before running L^{A}TEX the final time. Note, if you run BibTEX after modifying file.bbl, you will lose your modifications when L^{A}TEX is run again. Therefore, it is best to turn off write-permission on file.bbl after you have it correct.

• For a ‘tabular’ subfigure situation,^{10} the submit option may clip some of the subfigures. In fact, the way the submit option deals with subfigures needs work—to put it mildly.

Also, if your captions for each subfigure do not have the same number of lines when typeset, the subfigures will be misaligned. You can alleviate this by adding newlines (\\) to the shorter lines as necessary.

• Currently you have to make subfigure captions appear in bold font explicitly,

e.g.,

  \subfigure[\bf subcaption.]{\incfig{fig.eps}}

The next version of subfigure will remedy this.

• Using the starred versions of figure and table environments (floats which span both columns) inter-mixed with the unstarred versions (single-column wide floats) often creates a situation where the figures or tables appear out of order. For the final copy, this can normally be corrected with judicious use of float position specifiers.^{11} Unfortunately, this is stock, documented behavior for L^{A}TEX. A non-distributable solution exists, but no freely-available one has surfaced. A request has been filed with the TEXperts to remedy this situation.

• The notices come after the author footnotes. To produce the correct behavior requires significant changes to the \maketitle command.

• The simulated journal cover-pages do not have the correct font for the journal name—does anyone know where to get such a tall, bold, squashed helveticstyle font? It looks like Bitstream’s Aurora condensed…

10See sample documents smpaiaa.tex and smpsubf.tex.

11Recall, never use just [h], always give other options or that float might get ‘stuck’ and force it and all the following floats to the end of the document.

4 Acknowledgements

Bundling and documenting this aiaa distribution in docstrip format was done by using other packages as a model, particularly, Mats Dahlgren’s dropping [4] and Jeff Goldberg et al’s endfloat [6].

I want to thank the people of the comp.text.tex newsgroup, the TEX/L^{A}TEX Frequently Asked Questions maintainers, and various package authors for patiently answering my inane questions, in particular and in no particular order:

▷ Donald Arsenau (asnd@reg.triumf.ca)
▷ Robin Fairbairns (Robin.Fairbairns@cl.cam.ac.uk)

▷ Piet van Oostrum (piet@cs.ruu.nl)

▷ Jeroen Nijhof (nijhof@th.rug.nl)

▷ Steven Douglas Cochran (sdc+@cs.cmu.edu)

▷ Jeffrey Goldberg (J.Goldberg@cranfield.ac.uk)
▷ Mark Wooding (mdw@excessus.demon.co.uk)
▷ Paul Foley (mycroft@actrix.gen.nz)
▷ David Kastrup (dak@fsnif.neuroinformatik.ruhr-uni-bochum.de)
▷ Jerry Leichter (leichter@smarts.com)
▷ P. W. Daly (daly@linpwd.mpae.gwdg.de)
▷ David Carlisle (carlisle@goofy.zdv.Uni-Mainz.de)

5 Sending a Bug Report

The aiaa distribution is highly likely to contain bugs. Reports of bugs in the package are most welcome. However, I consider this to be a minimally “supported” package. I will do what I can, when I can—promising nothing. Before filing a bug report, please take the following actions:

1. Ensure your problem is not due to your own input file or package(s):

2. Ensure your problem is not covered in the section “Known Problems” above;

3. Try to locate the problem by writing a minimal L^{A}TEX input file which re-

produces the problem. Include the command

    \setcounter{errorcontextlines}{50}

in your input to provide extra context when things go awry;

4. Run your file through L^{A}TEX;

5. Send a description of your problem, the input file and the log file via e-mail

  to: w.l.kleb@larc.nasa.gov.

I am not in the business of answering generic TEX/L^{A}TEX questions; so if your problem appears to be such, I will let you know.

Enjoy(?) the everpresent deadline and enjoy your L^{A}TEX!—bil

References

[1] Michel Goossens, Frank Mittelbach, and Alexander Samarin. The L^{A}TEX Companion. Addison-Wesley, Reading, Massachusetts, 1994.

[2] Helmut Kopka and Patrick W. Daly. A Guide to L^{A}TEX 2ε: Document Prepartion for Beginners and Advanced Users. 2nd ed. Addison-Wesley, Reading, Massachusetts, 1994.

[3] Leslie Lamport. L^{A}TEX: A Document Preparation System. 2nd ed. Addison-Wesley, Reading, Massachusetts, 1994.

[4] Mats Dahlgren. dropping—A L^{A}TEX Macro for Dropping the First Character(s) of a Paragraph. June 1996. (version 0.1) Electronic Documentation.

[5] David Carlisle. Packages in the ‘graphics’ bundle. December 1995. Electronic Documentation.

[6] James Darrell McCauley and Jeff Goldberg. The endfloat Package. October 1995. (version 2.4i) Electronic Documentation.